8. [0/5 Points] DETAILS PREVIOUS ANSWERS LARLINALG8 7.1.021. Find the characteristic equation and the eigenvalues (and corres11. [-14 Points] DETAILS LARLINALG8 7.1.045. Consider the linear transformation T: R → Rh whose matrix A relative to the stanplease solve both as other did wrong plz
Answer #1

8 matix -2. a O 3 -2 0 2 a) The characteristic equation is JA-AI)20 2- 9 0 3-1 -2 -0 0 2-2 ne get Expand along wlumn I (2-d)1 0 - . = A - X = (-1, 1, D CM 4= 2. 2-2 b 0 - 2 O 0 3-2 N 0 1 0 0 0 0 2-2 To -) 0 다. 0 0 M 월드, 1320 - 192 0 11) 122 0 A 0 이we have matix 블 A- 4 -1 ៗ a) sind ure eigenvalues of A. TA-AI)20 4-1 1 7-1 2 (4-22 (7-2) +2=0 12- ||d+30 20 (2.5)(1-6)= 0 The-½ -/ 12 Bz 3 Jos - {}} B-U11) O:4:1 c) Now, BY here, number of eigenrecton (say, 2) is equal to the number of columns (say,